OpenStudy (anonymous):

Find the range and the first and third quartiles for the data set. 26, 27, 73, 78, 57, 46, 90

OpenStudy (anonymous):

Range = 90-26=64

OpenStudy (anonymous):

64..Range Is Subtracting The Highest Number ( 90 ) From The Lowest ( 26) 90-26 = 64